DIRC Operations Report

Wednesday 2007/12/26 to Wednesday 2008/01/02


Oncall: Jose Benitez (Rolf Andreassen on backup)
Fast Monitoring: Rolf Andreassen, Jose Benitez

Contents:

    PEP Activities     BaBar Activities     DIRC activities     Backgrounds     Scheduled activities     Pending Work     Links

Luminosities this week

PEP peak lumi: 7.5 x 1033 (last week: 5.5 x 1033)

PEP delivery: 2 fb-1 (last week: 1.84 fb-1)

BaBar integrated lumi:2 fb-1 (last week: 1.84 fb-1)

Run 7 total delivery: 5.41 fb-1

Run 7 total BaBar recorded: 5.41 fb-1 total, including 0 fb-1 off-peak data

PEP activities/issues this week

  • Vaccum problems are preventing the LER current from being raised above 1800mA.

    BaBar activities/issues this week

  • Relatively smooth running at the Y(3S).
  • 3:45 operations meetings are back after the break.

    DIRC activities/issues this week

  • Updated JAS references.
  • DFB 9 in sector 6 was found to be causing high data truncation errors. May be replaced next time the back doors are open.

    See DIRC logbook for details

    Lumi and Background this week